At constant pressure, what should be the percentage increase in the temperature in Kelvin for a 10% increase in volume -

Grade:11

1 Answers

Arun
6 years ago
Dear Ayush
 
V1/T1 = (V1 + 0.1* V1)/T2
Or
T2 = 1.1 T1
Increase in Temperature = 1.1 T1 - T1 = 0.1 T1
Hence 
% increase in temp = 0.1 T1* 100/T1 = 10 %
 
Regards,
Arun (AskIITians Forum Expert)